PDS_VERSION = PDS3 RECORD_TYPE = STREAM OBJECT = TEXT PUBLICATION_DATE = 1997-08-21 NOTE = "MGS RST Instrument Health Report HEA7233A.TXT" END_OBJECT = TEXT END This report is being submitted by Joe Twicken. Dick Simpson is back at Stanford today and will be submitting the Instrument Health report tomorrow. Trish Priest reported that MGS would be on the LGA yesterday between 00:00 and 00:58, 05:21 and 07:07, and 14:36 and 16:22. All times are ERT. These switches to LGA were confirmed by the Engineering data. Trish also reported that MGS would be on the LGA yesterday between 20:45 and 22:31. This switch was not confirmed by the Engineering status channels, but does seem to be confirmed by temperature changes in a number of Engineering channels. I do not know why there are no changes in status indicators to confirm the switch to the LGA. Engineering Data ================ Engineering Data Statistical Summary from CSV2STAT: First SCET: 1997-08-20T00:01:39.750 Last SCET: 1997-08-20T23:59:59.228 CHANNL CHANNEL TITLE DN-LO DN-HI EU-LO EU-HI EU-AVG/STDEV RECDS ------ ------------- ----- ----- --------- --------- ---------/-------- ----- F-0190 HGA_AZ_ANG 0 0 0.00E+00 0.00E+00 0.00E+00/0.00E+00 2084 F-0195 HGA_EL_ANG 0 0 0.00E+00 0.00E+00 0.00E+00/0.00E+00 2084 L-0034 MOT1_RANGING 0 0 17746 L-0036 MOT1_TLM_MOD 0 0 17746 L-0037 MOT1_USO_ENA 0 1 17746 L-0044 MOT2_RANGING 0 1 17746 L-0046 MOT2_TLM_MOD 0 1 17746 L-0047 MOT2_USO_ENA 0 1 17746 L-0053 RF_SW_OSC 2 2 17746 L-0060 TWTA1_FILMNT 0 1 17746 L-0061 TWTA1_HV 0 1 24535 L-0070 TWTA2_FILMNT 1 1 24535 L-0071 TWTA2_HV 0 1 17746 L-0080 USO_PWR 0 0 17746 L-0090 KaBLE_STAT 1 1 17746 L-0091 KaBLE_ENABLE 0 0 24535 L-0111 MOT1_RCV_AGC 206 255 -1.53E+02 -1.39E+02 -1.45E+02/4.71E-02 9984 L-0121 MOT2_RCV_AGC 155 255 -1.53E+02 -1.25E+02 -1.33E+02/8.83E-02 9984 L-0131 TWTA1_HLX_I 0 37 -3.84E-02 6.96E-01 2.23E-02/1.92E-03 9984 L-0132 TWTA1_ANOD_V 0 244 -1.00E+03 -5.17E+02 -1.00E+03/1.66E-01 8040 L-0141 TWTA2_HLX_I 17 43 2.86E-01 8.05E-01 7.29E-01/1.26E-03 9984 L-0142 TWTA2_ANOD_V 60 241 -8.68E+02 -5.06E+02 -8.57E+02/4.74E-01 8040 L-0200 USO_REG_V 203 205 4.06E+00 4.10E+00 4.08E+00/9.55E-05 8040 L-0201 USO_OVEN_V 123 126 2.46E+00 2.52E+00 2.48E+00/7.53E-05 8040 T-0302 EPC1_T 112 122 2.68E+01 3.11E+01 2.80E+01/1.55E-02 2216 T-0303 EPC2_T 98 117 2.90E+01 3.71E+01 3.48E+01/2.71E-02 2216 T-0310 RF_ISOLTR1_T 159 181 2.79E+01 3.56E+01 3.04E+01/1.66E-02 9983 T-0311 RF_ISOLTR2_T 167 185 2.64E+01 3.28E+01 3.11E+01/1.45E-02 9983 T-0312 TWT1_T 152 181 2.79E+01 3.82E+01 3.04E+01/3.38E-02 2216 T-0313 TWT2_T 160 189 2.49E+01 3.52E+01 3.37E+01/4.14E-02 2216 T-0315 USO_T 177 180 9.93E+00 1.09E+01 1.03E+01/3.76E-03 2216 T-0316 KA_AMP_T 136 154 3.74E+01 4.45E+01 4.02E+01/1.48E-02 4161 L-0200, L-0201 and T-0315 are nominal. Open Loop Data ============== Data Collection --------------- No new data. Closed Loop Data ================ Data Collection --------------- We have obtained the following files from OSCAR and EPIC: Original File Name Name on SOPC ------------------------------ ------------ ATDF: ODF: 97232C233.DAT;1 7232233A.ODF IONCAL: TROPCAL: WEATHER: EOP: